100ml of co2 was collected at 33degree C and 700 mm pressure. what will be its volume ,if the pressure is changed to 720mm at the same temperature?

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
PV / T = constant
 
P1 * V1 / T1  = P2 * V2 / T2
 
700 * 100 / 306 = 720 * V2 / 306
 
V2 = 700 * 100 / 720 = 97.22 ml
 
Regards
Arun (askIITians forum expert)